OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)). -7 -3 10 33

OpenStudy (anonymous):

First find f(g(x))

OpenStudy (anonymous):

Can you do that right now please?

OpenStudy (anonymous):

f(x)=x+8 x^2-6x-7+8 -13x+8 =x^2-5x

OpenStudy (anonymous):

\[f(g(x)) = (x^2-6x-7)+8\]

OpenStudy (anonymous):

\[f(g(2)) = ((2)^2-6(2)-7)+8\]

OpenStudy (anonymous):

(4-12-7)+8=-7

OpenStudy (anonymous):

yes